Assume you just purchased 100 shares of Home Depot at $20.46 per share, and 50 percent of this was purchased on the margin.
Answer #1

Qa) Total value = 20.46 × 100 = 2,046

Amount borrowed= 50% × 2,046 = - $1,023

Contribution = $1,023

Qb) Total value = 100 × 24.42 = $2,442

Loan= - $1,023

Margin= $1,419

Profit = 2,442 - 2,046 = $396